OpenStudy (anonymous):

4+1/x/3-1/x leave answer in factored form

OpenStudy (anonymous):

\[4 + \left.\left(\frac{1}{x}\right)\right/3 - \left(\frac{1}{x}\right)=4-\frac{2}{3 x}\]